OpenStudy (tteesstt8080):

Would anyone please help me find the value of these two expressions without using a calculator? cos(arccos 1 + arcsin 0) tan[arccos(-1)-arccos 1)]

OpenStudy (sshayer):

\[\cos \left( \cos^{-1} 1+\sin^{-1} 0 \right)=\cos \left( 0+0 \right)=\cos 0=1\]

OpenStudy (tteesstt8080):

Oh... wait but the inverse... the inverse confuses me...

OpenStudy (tteesstt8080):

so you just have to apply the unit circle? kind of like pretending inverse doesn't exist there?

OpenStudy (sshayer):

\[\tan [\cos^{-1} (-1)-\cos^{-1} 1]=\tan [ \pi-0]=\tan \pi=0\]

OpenStudy (amorfide):

\[\arccos(x)=\int\limits_{}^{} \frac{ - 1 }{ \sqrt{1-x^{2}} }\] let x=1 \[\arcsin(x)=\int\limits_{}^{} \frac{ 1 }{ \sqrt{1-x^{2}} }\] let x=0 then do cos of the sum

OpenStudy (tteesstt8080):

Oh okay! Thank you!

OpenStudy (tteesstt8080):

I think I understand now.